OpenStudy (anonymous):

write each equation in standard form using integers y= 3/4x +5 write each equation in standard form using integers y + 4 = 1/3 (x + 6)

OpenStudy (anonymous):

-(3 x)/4+y-5 = 0 -x/3+y-2 = 0